OpenStudy (anonymous):

Simplify. 5/(x^2 - xy) - 5/(y^2 - xy)

OpenStudy (anonymous):

First, you'll need to get same denominators: \[\frac{5}{x(x-y)} - \frac{5}{y(y-x)} = \frac{5(y)(y-x)-5(x)(x-y)}{-xy(x-y)}=\frac{-5y^2-5yx -5x^2 +5xy}{xy(x-y)(y-x)}\] \[= \frac{-5(y^2-x^2)}{xy(x-y)(y-x)}\] \[=\frac{-5(y-x)(y+x)}{xy(y-x)(x-y)} = \frac{-5(y+x)}{xy(x-y)} =\] Correct me if I'm wrong please ^_^

OpenStudy (anonymous):

oh remove the negative from the second step in the denominator ! >_< my bad lol

OpenStudy (anonymous):

and it's supposed to be 5y^2 and not -5y^2, sorry again ^_^"

OpenStudy (anonymous):

hope you got the point , sorry about the mess ^^"

OpenStudy (anonymous):

lol....I'm a bit confused now...but if it's only the negatives you messed up on I think I understand. Thanks!

OpenStudy (anonymous):

np :) lol just the second and third step, I added (-) out of the blue lol, you'll understand what I mean when you write it down on paper, silly mistake :)